# Mutual-visibility coloring of graphs

### v1

@@ 3 Mutual-visibility coloring in glued $t$-ary trees
Note that we have considered $1+(t-1)+t(t-1)+t^{2}(t-1)+\ldots+t^{i-2}(t-1)=1+t^{i-1}-1=t^{i-1}$ cycles. Since all the $t^{n}-1$ cycles $\chi^{n}_{m}$ are disjoint and since cardinality of elements in $\chi^{n}_{m}$ is at least 4i4n24i-4n-2, if $m\ncong 1\pmod{t}$ and 4i4n44i-4n-4, if $m\cong 1\pmod{t}$ implies that $|\bigcup_{m,n}\chi^{n}_{m}|\geq\sum_{n=1}^{i-1}\Bigl((t-2)t^{n-1}(4i-4n-2)+t^{n-1}(4i-4n-4)\Bigr)=\frac{2(t^{i}-1)}{t-1}$. Therefore, at least $\frac{2(t^{i}-1)}{t-1}$ colors are used three times to color $GT(r,t)$. But, we have only 2(ri)+12(r-i)+1 colors and $r\leq\frac{t^{i}-1}{t-1}+i-1$ so that $2(r-i)+1\leq\frac{2(t^{i}-1)}{t-1}-1$, a contradiction.
Therefore, at least 2(ri)+22(r-i)+2 colors are required for a mutual-visibility coloring of $GT(r,t)$. Now, let us investigate whether 2(ri)+22(r-i)+2 colors are sufficient. Suppose we have 2(ri)+22(r-i)+2 colors. Consider the cycles $C_{m}^{n}$, $n\in[i-2]$. Then by similar arguments, all these cycles are disjoint and cardinality of elements in $\chi^{n}_{m}$ is at least 4i4n44i-4n-4, if $m\ncong 1\pmod{t}$ and 4i4n64i-4n-6, if $m\cong 1\pmod{t}$, implies that $|\bigcup_{n,m}\chi^{m}_{n}|\geq\frac{2(t^{i-1}-1)}{t-1}+2t^{i-2}$, $n\in[i-2]$. Let $S$ be the collection of all vertices of depth (i1)(i-1) of $T_{r}^{(1)}$ that are not part of any $C_{m}^{n}$ for $n\in[i-2]$. Note that $|S|=t^{i-1}-2t^{i-2}$ and $|S|>0$ for $t\geq 3$. Let $T$ be the collection of quasi-leaves obtained by taking one quasi-leaf of the binary tree rooted at $k$, for each $k\in S$. Now consider the cycles $C_{k,l}$ where $k,l\in T$. In each of these cycles there are exactly 2(ri+3)2(r-i+3) vertices that are not part of any $C_{m}^{n}$, $n\in[i-2]$. Since we have only 2(ri)+22(r-i)+2 colors, at least two colors must be repeated at least three times in every cycle, and these colors are distinct. Using similar arguments, we can prove that these colors are also distinct from $\chi_{m}^{n}$, $n\in[i-2]$. Therefore, at least $\frac{2(t^{i-1}-1)}{t-1}-2t^{i-2}+t^{i-1}-2t^{i-2}=\frac{2(t^{i-1}-1)}{t-1}+t^{i-1}$ distinct colors must be used to color $GT(r,t)$. But in this case, if $r\leq\frac{t^{i-1}-1}{t-1}+\frac{t^{i-1}}{2}+i-2$, so that $2(r-i)+2\leq\frac{2(t^{i-1}-1)}{t-1}+t^{i-1}-2$. Hence, the theorem. ∎

### v2

@@ 3 Mutual-visibility coloring in glued $t$-ary trees
Note that we have considered $1+(t-1)+t(t-1)+t^{2}(t-1)+\ldots+t^{i-2}(t-1)=1+t^{i-1}-1=t^{i-1}$ cycles. Since all the $t^{n}-1$ cycles $\chi^{n}_{m}$ are disjoint and since cardinality of elements in $\chi^{n}_{m}$ is at least 4i4n24i-4n-2, if $m\not\equiv 1\pmod{t}$ and 4i4n44i-4n-4, if $m\equiv 1\pmod{t}$ implies that $|\bigcup_{m,n}\chi^{n}_{m}|\geq\sum_{n=1}^{i-1}\Bigl((t-2)t^{n-1}(4i-4n-2)+t^{n-1}(4i-4n-4)\Bigr)=\frac{2(t^{i}-1)}{t-1}$. Therefore, at least $\frac{2(t^{i}-1)}{t-1}$ colors are used three times to color $GT(r,t)$. But, we have only 2(ri)+12(r-i)+1 colors and $r\leq\frac{t^{i}-1}{t-1}+i-1$ so that $2(r-i)+1\leq\frac{2(t^{i}-1)}{t-1}-1$, a contradiction.
Therefore, at least 2(ri)+22(r-i)+2 colors are required for a mutual-visibility coloring of $GT(r,t)$. Now, let us investigate whether 2(ri)+22(r-i)+2 colors are sufficient. Suppose we have 2(ri)+22(r-i)+2 colors. Consider the cycles $C_{m}^{n}$, $n\in[i-2]$. Then by similar arguments, all these cycles are disjoint and cardinality of elements in $\chi^{n}_{m}$ is at least 4i4n44i-4n-4, if $m\not\equiv 1\pmod{t}$ and 4i4n64i-4n-6, if $m\equiv 1\pmod{t}$, implies that $|\bigcup_{n,m}\chi^{m}_{n}|\geq\frac{2(t^{i-1}-1)}{t-1}+2t^{i-2}$, $n\in[i-2]$. Let $S$ be the collection of all vertices of depth (i1)(i-1) of $T_{r}^{(1)}$ that are not part of any $C_{m}^{n}$ for $n\in[i-2]$. Note that $|S|=t^{i-1}-2t^{i-2}$ and $|S|>0$ for $t\geq 3$. Let $T$ be the collection of quasi-leaves obtained by taking one quasi-leaf of the binary tree rooted at $k$, for each $k\in S$. Now consider the cycles $C_{k,l}$ where $k,l\in T$. In each of these cycles there are exactly 2(ri+3)2(r-i+3) vertices that are not part of any $C_{m}^{n}$, $n\in[i-2]$. Since we have only 2(ri)+22(r-i)+2 colors, at least two colors must be repeated at least three times in every cycle, and these colors are distinct. Using similar arguments, we can prove that these colors are also distinct from $\chi_{m}^{n}$, $n\in[i-2]$. Therefore, at least $\frac{2(t^{i-1}-1)}{t-1}-2t^{i-2}+t^{i-1}-2t^{i-2}=\frac{2(t^{i-1}-1)}{t-1}+t^{i-1}$ distinct colors must be used to color $GT(r,t)$. But in this case, if $r\leq\frac{t^{i-1}-1}{t-1}+\frac{t^{i-1}}{2}+i-2$, so that $2(r-i)+2\leq\frac{2(t^{i-1}-1)}{t-1}+t^{i-1}-2$. Hence, the theorem. ∎